# Integrated Vision and Language Lab, KAIST, South Korea

### v1

@@ Abstract
This paper explores sentence-level Multilingual Visual Speech Recognition with a single model for the first time. As the massive multilingual modeling of visual data requires huge computational costs, we propose a novel strategy, processing with visual speech units. Motivated by the recent success of the audio speech unit, the proposed visual speech unit is obtained by discretizing the visual speech features extracted from the self-supervised visual speech model. To correctly capture multilingual visual speech, we first train the self-supervised visual speech model on 5,512 hours of multilingual audio-visual data. Through analysis, we verify that the visual speech units mainly contain viseme information while suppressing non-linguistic information. By using the visual speech units as the inputs of our system, we pre-train the model to predict corresponding text outputs on massive multilingual data constructed by merging several VSR databases. As both the inputs and outputs are discrete, we can greatly improve the training efficiency compared to the standard VSR training. Specifically, the input data size is reduced to 0.016% of the original video inputs. In order to complement the insufficient visual information in speech recognition, we apply curriculum learning where the inputs of the system begin with audio-visual speech units and gradually change to visual speech units. After pre-training, the model is finetuned on continuous features. We set new state-of-the-art multilingual VSR performances by achieving comparable performances to the previous language-specific VSR models, with a single trained model.
Index Terms—  Visual Speech Recognition, Lip Reading, Visual Speech Units, Discrete Speech Units, Multilingual VSR
1 Introduction
@@
These days, speech processing technologies have made great progress in diverse applications such as speech recognition [1, 2, 3, 4], speech synthesis [5, 6, 7, 8, 9], and speech translation [10, 11, 12, 13]. Now, it is easy to find a speech processing model that can proficiently handle approximately 100 languages [14, 15]. However, multilingualism has been mainly explored for audio-based speech processing systems [16, 17], while visual-based speech processing systems are still tied to developing monolingual systems [18, 19, 20]. There are two reasons for the lagging development of visual-based speech processing systems: 1) The high dimensionality of visual data compared to audio puts a challenge in training a large-scale model with massive multilingual data. Compared to the same length of audio, visual data requires about six times larger bits [21] in a standard visual speech recognition process [22]. Moreover, the requirement of encoding spatial information using two-dimensional convolutions also increases the computation costs of visual speech processing compared to its counterpart. 2) The low quantity of labeled data in visual speech processing systems presents a formidable obstacle to technology development. In contrast to the tremendous amount of publicly available audio-text data [23], a very limited number of video-text data are available, especially for non-English [19].
@@
Fig. 1: Illustration of the proposed multilingual VSR framework. (a) Visual speech units are obtained by quantizing the visual speech features extracted from mAV-HuBERT. (b) The Transformer encoder-decoder model is pre-trained with discrete inputs and outputs. The task difficulty is gradually increased by using progressive masking $\mathcal{M}$. Pre-training commences with audio-visual speech units as inputs, and these inputs gradually transition to visual speech units through progressive masking of the audio speech units. (c) After pre-training the model with discrete inputs and outputs, it is finetuned with continuous features to boost the VSR performances.
@@
Specifically, we introduce multilingual Audio-Visual Hidden Unit BERT (mAV-HuBERT) which is the self-supervised speech model for extracting the visual speech unit. We train the mAV-HuBERT on 5,512 hours of multilingual audio-visual data composed of nine languages. Therefore, we can correctly put the multilingual viseme information into our visual speech unit. Then, we pre-train an encoder-decoder model by setting the inputs with the visual speech units and the outputs with the corresponding text, forming a unit-to-unit translation framework (i.e., translation between discrete tokens) [13]. Inspired by the recent successes of VSR that leverage audio modal information to complement limited visual data [34, 35, 36, 37, 38], we propose to use curriculum learning with a gradual increase in task difficulty using audio modality. Concretely, our unit-to-unit pre-training is initiated with audio-visual inputs and then gradually changed to visual inputs. With this curriculum learning, the model can find the optimization points stable and achieve higher performance with the complemented information of multi-modality. To mitigate the comparatively small amount of public visual-text paired data, we utilize the recently proposed automatic labels of [39] and [20] where the text labels are obtained by their automatic labeling processes. Finally, the pre-trained model is finetuned with continuous input features to maximize the VSR performances.
The major contributions of this paper can be summarized as follows: 1) To the best of our knowledge, this is the first work exploring sentence-level multilingual VSR with a single trained model. 2) We propose to use a visual speech unit which is a quantized visual speech feature extracted from a self-supervised model, in modeling visual speech representations. 3) In order to correctly capture the visemes from multilingual visual speech, we train an audio-visual self-supervised speech model on 5,512 hours of multilingual data. 4) By employing the visual speech unit, we can train the multilingual VSR model with discrete inputs and outputs (i.e., text). With this, we can drastically reduce the computational costs and accelerate the pre-training time by about 10 times compared to the standard training. 5) Through analysis, we verify that the visual speech unit mainly holds the viseme information while suppressing non-linguistic variations, enabling VSR training with discrete inputs. We set new state-of-the-art VSR performances by achieving comparable performances with the multiple previous monolingual VSR methods.
2 Related Work
2.1 Visual Speech Recognition (VSR)
@@
VSR aims to predict the spoken words from silent lip movements video. Early works [40, 41, 42, 43] focused on word-level VSR by using CNN [44] and the RNN [45, 46]. Large-scale lip-reading sentence datasets [47, 48] have boosted the development of sentence-level VSR. By employing Transformer [49] architecture, [50] proposed a powerful sentence-level end-to-end VSR model. Moreover, the integration of the hybrid CTC/Attention objective [51, 52] into VSR, greatly improved the recognition performances. Recent VSR technologies [22, 53, 54, 39] also employed transformer-variant architectures and improved the VSR performances. For advanced training strategies, many researchers try to reduce the gap between visual and audio modalities. They [34, 55, 35, 27, 36, 56, 57] studied how to effectively transfer audio knowledge into the VSR model by using knowledge distillation [58] and memory network [59]. However, these previous VSR approaches have mainly developed for high-resource languages, English and Mandarin [60]. VSR for different languages, especially low VSR resource languages, has only been addressed recently [18, 61, 19, 20]. In particular, a recent approach [20] proposed the labeled data for low VSR resource languages using automatic labeling processes.
@@
This paper is the first work exploring sentence-level multilingual VSR with a single model. To mitigate the huge computational costs in training the multilingual VSR model, we propose to pre-train the model with discrete inputs and outputs by using visual speech units. To complement the low amount of video-text data, we bring the automatic labels of [39, 20] and propose curriculum learning that utilizes audio modality to provide rich speech information.
2.2 Audio speech unit
@@
Speech unit [29] is the discretized speech representation of self-supervised speech models such as HuBERT [31], Wav2Vec2.0 [30], and WavLM [62]. It is possible to suppress non-linguistic features and mainly keep the linguistic contents by selecting proper layers to extract the speech features [29, 63]. By using the speech unit as pseudo text, Textless Natural Language Processing becomes possible [12, 64, 65, 13, 66]. Moreover, speech units have promising potential to be used in multi-modal processing as they greatly reduce the data size [33, 67, 21].
@@
Table 1: Dataset statistics for pre-training mAV-HuBERT and training multilingual VSR model. Audio-visual data is utilized for training mAV-HuBERT, while video-text data is utilized for training the multilingual VSR model (i.e., for both pre-training and finetuning).
@@
Motivated by this, we propose a visual speech unit which is the quantized representation of visual speech features extracted from the self-supervised visual speech model [37]. We introduce mAV-HuBERT for multilingual visual speech unit extraction. We show that the proposed visual speech unit contains mainly viseme information while suppressing the other characteristics.
3 Method
3.1 Multilingual Visual Speech Unit Extraction
Audio speech units [29] can be obtained by clustering the speech features of a self-supervised speech model such as HuBERT [31]. Analogous to audio speech units, we propose visual speech units, which can be obtained by quantizing the visual speech features derived from a pre-trained visual speech model. In order to get visual speech units, we choose AV-HuBERT [37] for the self-supervised visual speech model, which is well-known for its discriminative visual speech features. However, AV-HuBERT is pre-trained on English-only audio-visual data, which deviates from our primary objective of achieving multilingualism. Hence, we initially introduce a multilingual variant of AV-HuBERT (mAV-HuBERT) to ensure the accurate incorporation of multilingual viseme information into the final visual speech units. To this end, we train the model on 5,512 hours of multilingual dataset composed of 9 languages (En, Es, It, Fr, Pt, De, Ru, Ar, El) by merging LRS2 [47], LRS3 [48], VoxCeleb2 [68], and AVSpeech [69]. As VoxCeleb2 and AVSpeech do not have language identities, we obtain the language identity of each utterance by using a pre-trained language identifier of Whisper [15], to select the data. For the prediction target of the masked prediction of mAV-HuBERT, we use clusters of speech features obtained from a pre-trained multilingual HuBERT [31, 12]. We use the target size of 1,000 and train the model for 350k steps with one iteration. We show that our mAV-HuBERT is more suitable for multilingual speech modeling than the previous English-trained AV-HuBERT in Sec. 4.3.1.
With the pre-trained mAV-HuBERT, we extract the visual speech unit by clustering (i.e., quantizing) the output visual speech features, as shown in Fig. 1(a). Please note that AV-HuBERT can extract both audio-only features and visual-only features through its modality dropout. Hence, we only use the visual inputs to extract the visual speech units. For the token size of the visual speech unit, we use 1,000, so each visual speech unit can be represented with 10 bits. Please note that one video frame having grayscale and 88 $\times$ 88 size (i.e., the standard for VSR) requires 61,952 bits [22, 37]. Therefore, we can reduce the data size to 0.016% compared to the raw visual inputs, which enables us to greatly increase the training batch size and reduce training speed by removing the visual front-end (e.g., 2D CNNs). We analyze the efficiency of visual speech units by comparing them with the standard raw inputs in Sec. 4.3.2.
3.2 Pre-training: Visual Speech Unit to Text Translation
@@
By setting the inputs with visual speech units, we pre-train our model to predict corresponding text. Therefore, now the inputs and outputs are both discrete, which is illustrated in Fig. 1(b). As the visual speech units mainly contain linguistic information, we can pre-train the model to construct the knowledge of visual speech modeling even by using discrete inputs. We analyze which information is mainly held by the visual speech units in Sec. 4.3.3.
@@
Nevertheless, translating visual speech units directly into output text from scratch is challenging for the model in identifying optimal solutions. Since visual information contains scarce speech information [34, 36, 35] compared to audio, training the model directly to perform visual-to-text conversion might be hard to find the solution. To mitigate this, we bring the motivation from the recent success of VSR, which utilizes auxiliary audio information during training [55, 34, 37, 35, 56, 18, 57]. Specifically, we initiate the pre-training with audio-visual speech units where both audio speech units and visual speech units are utilized as inputs, similar to [37, 70]. Then, we gradually masked out the audio part as the pre-training progressed, resulting in the final training stage exclusively utilizing visual speech units as inputs. Therefore, the model can easily find the optimization points through this curriculum learning, with the aid of complementary audio information. Concretely, the embeddings of the visual speech units $f_{v}\in\mathbb{R}^{T\times D}$ and audio speech units $f_{a}\in\mathbb{R}^{T\times D}$ are concatenated as, $g=\mathcal{M}(f_{a})\oplus f_{v}$, where $T$ is the sequence length, $D$ is the dimension of embedding, $\oplus$ represents concatenation operation in the embedding dimension, and $g\in\mathbb{R}^{T\times 2D}$ is the concatenated feature. $\mathcal{M}(\cdot)$ is a masking function that randomly masks out $p\%$ of frames from the input sequence. We progressively increase the $p$ from 0 to 100 as the pre-training progresses, so the transition from audio-visual inputs to visual inputs can be made. The effectiveness of the curriculum learning using input transition from audio-visual to visual can be found in Sec. 4.3.4.
@@
Then, we reduce the embedding dimension of the concatenated feature $g$ using a linear layer. Here, we provide the language information by providing the language embedding which will be added to the feature, following [71]. Finally, through the Transformer encoder-decoder architecture, we translate the visual inputs into the output text in an auto-regressive manner. The objective function of our learning problem can be represented as follows,
$\mathcal{L}=-\sum_{s=1}^{S}\log P(y_{s}|\mathbf{X},y_{<s}),$ (1)
@@
where $y_{s}$ is the text annotation for current step $s$ and $y_{<s}$ is the previous outputs, $\mathbf{X}$ is the input speech units, and $S$ is the length of the text. For multilingual training, we use five languages (En, Pt, Es, Fr, It) by merging LRS3 [48], mTEDx [72], automatic labels for En of [39], and automatic labels for Pt, Es, Fr, and It of [20], forming 4,545 hours of data. We summarize the dataset statistics in Table 1.
3.3 Finetuning: Multilingual Visual Speech Recognition

### v2

# Efficient Training for Multilingual Visual Speech Recognition: Pre-training with Discretized Visual Speech Representation

@@ Abstract.
This paper explores sentence-level multilingual Visual Speech Recognition (VSR) that can recognize different languages with a single trained model. As the massive multilingual modeling of visual data requires huge computational costs, we propose a novel training strategy, processing with visual speech units. Motivated by the recent success of the audio speech unit, we propose to use a visual speech unit that can be obtained by discretizing the visual speech features extracted from the self-supervised visual speech model. Through analysis, we verify that the visual speech units mainly contain viseme information while suppressing non-linguistic information. By using the visual speech units as the inputs of our system, we propose to pre-train a VSR model to predict corresponding text outputs on multilingual data constructed by merging several VSR databases. As both the inputs (i.e., visual speech units) and outputs (i.e., text) are discrete, we can greatly improve the training efficiency compared to the standard VSR training. Specifically, the input data size is reduced to 0.016% of the original video inputs. In order to complement the insufficient visual information in speech recognition, we apply curriculum learning where the inputs of the system begin with audio-visual speech units and gradually change to visual speech units. After pre-training, the model is finetuned on continuous features. We set new state-of-the-art multilingual VSR performances by achieving comparable performances to the previous language-specific VSR models, with a single trained model.
Visual Speech Recognition, Lip Reading, Multilingual VSR
Both authors have contributed equally to this work. Corresponding Author.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 28 - November 1, 2024; Melbourne, Australia.isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
@@
These days, speech processing technologies have made great progress in diverse applications such as speech recognition (Guo et al., 2021; Shim et al., 2021; Prabhavalkar et al., 2023; Hong et al., 2023), speech synthesis (Wu et al., 2022; Choi et al., 2023; Zhang et al., 2023; Jiang et al., 2023; Maiti et al., 2023), and speech translation (Inaguma et al., 2019; Jia et al., 2022; Lee et al., 2022; Kim et al., 2023a). Now, it is easy to find a speech processing model that can proficiently handle approximately 100 languages (Adams et al., 2019; Radford et al., 2023). However, multilingualism has been mainly explored for audio-based speech processing systems (Toshniwal et al., 2018; Lux et al., 2022), while visual-based speech processing systems are still tied to developing monolingual systems (Ma et al., 2022a; Kim et al., 2023c; Yeo et al., 2023b). There are two reasons for the lagging development of visual-based speech processing systems: 1) The high dimensionality of visual data compared to audio puts a challenge in training a large-scale model with massive multilingual data. Compared to the same length of audio, visual data requires about six times larger bits (Kim et al., 2023b) in a standard visual speech recognition process (Ma et al., 2021b). Moreover, the requirement of encoding spatial information using two-dimensional convolutions also increases the computation costs of visual speech processing compared to its counterpart. 2) The low quantity of labeled data in visual speech processing systems presents a formidable obstacle to technology development. In contrast to the tremendous amount of publicly available audio-text data (Pratap et al., 2020), a very limited number of video-text data are available, especially for non-English (Kim et al., 2023c).
@@
Specifically, we employ AV-HuBERT (Shi et al., 2021), a self-supervised visual speech model, to extract our visual speech units. We newly train the AV-HuBERT model on 5,512 hours of multilingual audio-visual data composed of nine languages, given that the original AV-HuBERT was trained solely on English audio-visual data. To differentiate between the English-only trained and multilingual-trained AV-HuBERT models, we denote the latter as mAV-HuBERT. With the mAV-HuBERT, we can correctly capture the multilingual viseme information into our visual speech unit. Then, we propose to pre-train an encoder-decoder model by setting its inputs with the visual speech units and the outputs with the corresponding text, forming a unit-to-unit translation framework (i.e., translation between discrete tokens) (Kim et al., 2023a). Moreover, inspired by the recent successes of VSR that leverage audio modal information to complement limited visual data (Zhao et al., 2020; Ren et al., 2021; Kim et al., 2021; Shi et al., 2021; Haliassos et al., 2022), we propose to use curriculum learning with a gradual increase in task difficulty using audio modality. Concretely, our unit-to-unit pre-training is initiated with audio-visual inputs and then gradually changed to visual inputs. With this curriculum learning, the model can find the optimization points stable and achieve higher performance with the complementary multi-modal information. To mitigate the comparatively small amount of public visual-text paired data, we utilize the recently proposed automatic labels of (Ma et al., 2023) and (Yeo et al., 2023b) where the text labels are obtained by their automatic labeling processes. Finally, the pre-trained model is finetuned with continuous input features to maximize the VSR performances.
Figure 1. Illustration of the proposed multilingual VSR framework. (a) Visual speech units are obtained by quantizing the visual speech features extracted from mAV-HuBERT. (b) The Transformer encoder-decoder model is pre-trained with discrete inputs and outputs. The task difficulty is gradually increased by using progressive masking $\mathcal{M}$. Pre-training commences with audio-visual speech units as inputs, and these inputs gradually transition to visual speech units through progressive masking of the audio speech units. (c) After pre-training the model with discrete inputs and outputs, it is finetuned with continuous features to boost the VSR performances.
The major contributions can be summarized as follows:
  • To the best of our knowledge, this is the first work exploring sentence-level multilingual VSR with a single trained model. The proposed VSR model can recognize five different languages using a single model, whereas previous methods required training separate models for each language.
  • We propose to employ visual speech units as inputs to pre-train the multilingual VSR model, thereby establishing discrete inputs and outputs (i.e., text). With this, we can drastically reduce the computational costs and accelerate the pre-training time by about 10 times compared to the standard training.
  • Through analysis, we verify that the visual speech unit mainly holds the viseme information while suppressing non-linguistic features, enabling VSR training even by employing discretized inputs.
  • We set new state-of-the-art multilingual VSR performances by achieving comparable performances with the multiple previous monolingual VSR methods.
2. Related Work
2.1. Visual Speech Recognition (VSR)
@@
Visual Speech Recognition (VSR) aims to predict the spoken words from silent lip movements video. Early works (Chung and Zisserman, 2017b; Stafylakis and Tzimiropoulos, 2017; Petridis et al., 2017, 2018a) focused on word-level VSR by using CNN (He et al., 2016) and the RNN (Chung et al., 2014; Hochreiter and Schmidhuber, 1997). Large-scale lip-reading sentence datasets (Chung et al., 2017; Afouras et al., 2018a) have boosted the development of sentence-level VSR. By employing Transformer (Vaswani et al., 2017) architecture, (Afouras et al., 2018b) proposed a powerful sentence-level end-to-end VSR model. Moreover, the integration of the hybrid CTC/Attention objective (Watanabe et al., 2017; Petridis et al., 2018b) into VSR, greatly improved the recognition performances. Recent VSR technologies (Ma et al., 2021b; Prajwal et al., 2022; Chang et al., 2023a; Ma et al., 2023) also employed transformer-variant architectures and improved the VSR performances. For advanced training strategies, many researchers try to reduce the gap between visual and audio modalities. They (Zhao et al., 2020; Afouras et al., 2020; Ren et al., 2021; Ma et al., 2021a; Kim et al., 2021, 2022; Yeo et al., 2023a) studied how to effectively transfer audio knowledge into the VSR model by using knowledge distillation (Hinton et al., 2015) and memory network (Weston et al., 2015). However, these previous VSR approaches have mainly developed for high-resource languages, English and Mandarin (Luo et al., 2020). VSR for different languages, especially low VSR resource languages, has only been addressed recently (Ma et al., 2022a; Zinonos et al., 2023; Kim et al., 2023c; Yeo et al., 2023b). In particular, a recent approach (Yeo et al., 2023b) proposed the labeled data for low VSR resource languages using automatic labeling processes.
@@
This paper is the first work exploring sentence-level multilingual VSR with a single model. To mitigate the huge computational costs in training the multilingual VSR model, we propose to pre-train the model with discrete inputs and outputs by using visual speech units. To complement the low amount of video-text data, we bring the automatic labels of (Ma et al., 2023; Yeo et al., 2023b) and propose curriculum learning that utilizes audio modality to provide rich speech information.
2.2. Audio speech unit
@@
Audio speech unit (Lakhotia et al., 2021) is the discretized speech representation of self-supervised speech models such as HuBERT (Hsu et al., 2021), Wav2Vec2.0 (Baevski et al., 2020), and WavLM (Chen et al., 2022). It is possible to suppress non-linguistic features and mainly keep the linguistic contents by selecting proper layers to extract the speech features (Lakhotia et al., 2021; Polyak et al., 2021). By using the speech unit as pseudo text, Textless Natural Language Processing becomes possible (Lee et al., 2022; Huang et al., 2022; Popuri et al., 2022; Kim et al., 2023a; Nguyen et al., 2023). Moreover, speech units have promising potential to be used in multi-modal processing as they greatly reduce the data size (Chang et al., 2023b; Park et al., 2023; Kim et al., 2023b).
Motivated by this, we propose to employ a visual speech unit which is the quantized representation of visual speech features extracted from the self-supervised visual speech model (Shi et al., 2021). We analyze the characteristics of visual speech units and show that the visual speech unit contains mainly viseme information while suppressing the other characteristics.
3. Method
3.1. Multilingual Visual Speech Unit Extraction
Audio speech units (Lakhotia et al., 2021) can be obtained by clustering the speech features of a self-supervised speech model such as HuBERT (Hsu et al., 2021). Analogous to audio speech units, we propose to employ visual speech units, which can be obtained by quantizing the visual speech features derived from a pre-trained visual speech model. In order to get visual speech units, we choose AV-HuBERT (Shi et al., 2021) for the self-supervised visual speech model, which is well-known for its discriminative visual speech features. However, AV-HuBERT is pre-trained on English-only audio-visual data, which deviates from our primary objective of achieving multilingualism. Hence, we initially train a multilingual variant of AV-HuBERT (mAV-HuBERT) to ensure the accurate incorporation of multilingual viseme information into the final visual speech units. To this end, we train the model on 5,512 hours of multilingual dataset composed of 9 languages (En, Es, It, Fr, Pt, De, Ru, Ar, El) by merging LRS2 (Chung et al., 2017), LRS3 (Afouras et al., 2018a), VoxCeleb2 (Chung et al., 2018), and AVSpeech (Ephrat et al., 2018). As VoxCeleb2 and AVSpeech do not have language identities, we obtain the language identity of each utterance by using a pre-trained language identifier of Whisper (Radford et al., 2023), to select the data. For the prediction target of the masked prediction of mAV-HuBERT, we use clusters of speech features obtained from a pre-trained multilingual HuBERT (Hsu et al., 2021; Lee et al., 2022). We use the target size of 1,000 and train the model for 350k steps with one iteration. Through analysis we confirm that mAV-HuBERT is more suitable for multilingual speech modeling than the English-only trained AV-HuBERT in Sec. 4.3.1.
With the pre-trained mAV-HuBERT, we extract the visual speech unit by clustering (i.e., quantizing) the output visual speech features, as shown in Fig. 1(a). Please note that AV-HuBERT can extract both audio-only features and visual-only features through its modality dropout. Hence, we only use the visual inputs to extract the visual speech units. For the token size of the visual speech unit, we use 1,000 so that each visual speech unit can be represented with just 10 bits. Please note that one video frame having grayscale and 88 $\times$ 88 size (i.e., the standard for visual speech recognition) requires 61,952 bits (Ma et al., 2021b; Shi et al., 2021). Therefore, we can reduce the data size to 0.016% compared to the raw visual inputs, which enables us to greatly increase the training batch size and accelerate the training speed by removing the visual front-end (e.g., 2D CNNs). We analyze the efficiency of visual speech units by comparing them with the standard raw inputs in Sec. 4.3.2.
@@ 3.1. Multilingual Visual Speech Unit Extraction
Table 1. Dataset used for pre-training mAV-HuBERT and training the multilingual VSR model. Audio-visual data is utilized for mAV-HuBERT, while video-text data is utilized for the multilingual VSR model (i.e., for both pre-training and finetuning).
@@
3.2. Pre-training: Visual Speech Unit to Text Translation
@@
By representing all training video data into visual speech units, we can significantly reduce the data size, enabling efficient training of a VSR model on large-scale multilingual data. Based on this key concept, we propose to pre-train our model to predict text by setting the inputs with visual speech units. Therefore, now the inputs and outputs are both discrete, which is illustrated in Fig. 1(b). As the visual speech units mainly contain linguistic information, we can pre-train the model to construct the knowledge of visual speech modeling even by using discrete inputs. We analyze the information contained in the visual speech units and validate how it can be worked, in Sec. 4.3.3.
@@
Nevertheless, translating visual speech units directly into output text from scratch is challenging for the model in identifying optimal solutions. Since visual information contains scarce speech information (Zhao et al., 2020; Kim et al., 2021; Ren et al., 2021) compared to audio, training the model directly to perform visual-to-text conversion might be hard to find the solution. To mitigate this, we bring the motivation from the recent success of VSR, which utilizes auxiliary audio information during training (Afouras et al., 2020; Zhao et al., 2020; Shi et al., 2021; Ren et al., 2021; Kim et al., 2022; Ma et al., 2022a; Yeo et al., 2023a). Specifically, we initiate the pre-training with audio-visual speech units where both audio speech units and visual speech units are utilized as inputs, similar to (Shi et al., 2021; Djilali et al., 2023). Then, we gradually masked out the audio speech units as the pre-training progressed, resulting in the final training stage exclusively utilizing visual speech units as inputs. Therefore, the model can easily find the optimization points through this curriculum learning, with the aid of complementary audio speech information. Concretely, the embeddings of the visual speech units $f_{v}\in\mathbb{R}^{T\times D}$ and audio speech units $f_{a}\in\mathbb{R}^{T\times D}$ are concatenated as, $g=\mathcal{M}(f_{a})\oplus f_{v}$, where $T$ is the sequence length, $D$ is the dimension of embedding, $\oplus$ represents concatenation operation in the embedding dimension, and $g\in\mathbb{R}^{T\times 2D}$ is the concatenated feature. $\mathcal{M}(\cdot)$ is a masking function that randomly masks out $p\%$ of frames from the input sequence. We progressively increase the $p$ from 0 to 100 as the pre-training progresses so that the transition from audio-visual inputs to visual inputs can be made. The effectiveness of this curriculum learning using input transition from audio-visual to visual can be found in Sec. 4.3.4.
@@
Then, we reduce the embedding dimension of the concatenated feature $g$ using a linear layer. Here, we provide the language information by providing the language embedding which will be added to the feature, following (Conneau and Lample, 2019). Finally, through the Transformer encoder-decoder architecture, we translate the visual inputs into the output text in an auto-regressive manner. The objective function of our learning problem can be represented as follows,
(1) $\mathcal{L}=-\sum_{s=1}^{S}\log P(y_{s}|\mathbf{X},y_{<s}),$
@@
where $y_{s}$ is the text annotation for current step $s$ and $y_{<s}$ is the previous outputs, $\mathbf{X}$ is the input speech units, and $S$ is the length of the text. For multilingual training, we use five languages (En, Pt, Es, Fr, It) by merging LRS3 (Afouras et al., 2018a), mTEDx (Elizabeth et al., 2021), automatic labels for En of (Ma et al., 2023), and automatic labels for Pt, Es, Fr, and It of (Yeo et al., 2023b), forming 4,545 hours of data. We summarize the dataset statistics in Table 1.
3.3. Finetuning: Multilingual Visual Speech Recognition